# Fact-Checking with Contextual Narratives: Leveraging Retrieval-Augmented LLMs for Social Media Analysis

### v1

@@ 3.4. Explainable judgment
Our Judgment Phase comprises key components of dealing with noisy evidence, comparison with the Claim leading to an interpretable explanation and coherent verdict.
3.4.1. Dealing with noisy evidence
Web-scraped evidence is often noisy and may not directly relate to the claim. This issue is exacerbated in older datasets, where images are frequently repurposed in different contexts across platforms. For verbose social media posts, retrieving relevant evidence relies on extracting factual search terms from claims. Additionally, if a claim’s images and text do not originate from published news articles, retrieval becomes more challenging since such content is rarely indexed by search engines. This challenge is further aggravated when the post is not in English or does not concern prominent Western contexts. Prior work highlights these limitations, such as for Spanish tweets (Dey et al., 2024) and image contextualization in East Africa and South Asia (Tonglet et al., 2025). To handle noisy evidence, we employ two main strategies:
Visual threshold-based filtering
We filter evidence using an image similarity threshold, following prior approaches (Tonglet et al., 2025; Dey et al., 2024). Our visual features are extracted from pre-trained networks that capture key visual aspects: faces, scenes, and visual semantics. Faces are detected using (Zhang et al., 2016), with the most prominent face embedded via (Schroff et al., 2015), yielding $\mathcal{V}^{face}\in\mathbb{R}^{512}$. Scene information is extracted via a model trained on 365 place categories (Zhou et al., 2017), producing $\mathcal{V}^{place}\in\mathbb{R}^{2048}$. Visual semantics are captured using (Dosovitskiy et al., 2020), resulting in $\mathcal{V}^{sem}\in\mathbb{R}^{1000}$. The final visual representation of an image is:
$\mathcal{V}=[\mathcal{V}^{face},\mathcal{V}^{place},\mathcal{V}^{sem}].$
The cosine similarity between $\mathcal{V}^{I_{P}}$ and $\mathcal{V}^{I_{k}}$ defines the final similarity score of an evidence image $I_{k}$ $\epsilon$ $\mathcal{C}_{k}^{\prime}$ with the claim image $I_{P}$. Following (Tonglet et al., 2025) we leverage the large training set of NewsCLIPpings to determine our similarity threshold of 0.9. Since reverse image searches are designed to retrieve high-quality, contextually relevant images, evidence with low visual similarity—typically originating from text-based searches—can be safely disregarded, as it is unlikely to offer meaningful context for verification.
Text-Based narrative clustering
While low visual similarity reliably indicates irrelevant evidence, text similarity alone is not a sufficient filtering criterion. Visually similar yet textually different evidence may indicate out-of-context usage rather than irrelevance. However texts with different narratives coupled with irrelevant texts—whether due to scraping failures or website restrictions—can introduce noise during decision-making. To address this, we use narrative clustering to group texts about related topics into similar clusters. This helps maintain distinct narratives while preventing conflicted narratives from being mixed.
By using visual similarity for filtering and clustering noisy textual evidence, we ensure that relevant evidence, even if contextually different, is retained and presented in a structured manner to the LLM for decision-making.
3.4.2. Prompt-based comparison and explanation
@@
Our prompt, Fig.3, evaluates the veracity of a social media post $P=(T_{P},I_{P})$, consisting of a textual claim $T_{P}$ and an associated image $I_{P}$, by leveraging clustered evidence obtained from both direct and reverse image search. Each narrative cluster $\mathcal{C}_{k}^{\prime}$ contains evidence from text-based search ($\mathcal{X}_{T_{k}}$) and evidence from reverse search ($\mathcal{X}_{I_{k}}$).
Narrative assessment
@@
The judgment process begins with narrative assessment, where the LLM compares the textual content of each evidence item in $\mathcal{C}_{k}^{\prime}$ against the claim text $T_{P}$. Unlike prior methods that treat context prediction as a standalone step (Tonglet et al., 2025), our method integrates contextual understanding directly into a unified prompting strategy. Using a 5W1H-inspired questioning framework—focusing on who, what, when, where, why, and how—the LLM evaluates alignment between the claim and evidence across five critical dimensions: named entities, dates, locations, key topics, and common objects. This entity-level comparison helps determine whether an evidence item supports the claim, conflicts with it, or lacks relevant content entirely. In real-world scenarios, each narrative cluster often includes a mixture of evidence with varying degrees of fine-grained support. Such variability introduces noise into the reasoning process, making it challenging to draw accurate conclusions if relying solely on the highest-ranked evidence within a narrative. To address this, we aggregate evidence holistically at the cluster level, treating each $\mathcal{C}_{k}^{\prime}$ as a coherent narrative. Rather than favoring individual pieces of evidence in isolation, the LLM considers the collective alignment pattern within each cluster. The output of this stage is a structured set of alignment labels for each narrative, guiding further reasoning.
Reasoning with narratives
@@
Following narrative assessment, where each narrative cluster $\mathcal{C}_{k}^{\prime}$ is labeled based on entity-level alignment, the LLM proceeds to reason over these narratives to derive a final veracity decision. This reasoning process is grounded in the principle that narrative coherence, especially when supported by image provenance, is critical to verifying factual claims. Specifically, the LLM first identifies whether there exists any narrative cluster that both (i) contains valid reverse image search evidence ($\mathcal{X}_{I_{k}}$), and (ii) has been labeled as Supports during narrative assessment. If such a narrative is found, the post is immediately labeled as True. If no supporting narrative with valid reverse evidence is available, the model then checks for supporting narratives that contain only text-based evidence ($\mathcal{X}_{T_{k}}$). If at least one such narrative supports the claim, the label remains True. However, if all narratives—whether supported by image or text searches—are either labeled as Conflicts or Irrelevant, the post is ultimately labeled as Misleading. This structured approach ensures that multiple sources of evidence are considered in a tiered fashion, reducing the risk of false negatives or misclassifications due to noisy or incomplete top-ranked results. Finally, the LLM produces a natural language explanation grounded in the alignment rationale, clearly articulating the justification behind the final verdict and enhancing the transparency and interpretability of the model’s reasoning.
LLM-Based veracity reasoning Inputs: Claim text: $T_{P}$ Clustered evidence: $\left\{\mathcal{C}_{k}^{\prime}\right\}_{k=1}^{K}$, where each cluster $\mathcal{C}_{k}^{\prime}$ contains: Direct search evidence: $\mathcal{X}_{T_{k}}$ Reverse image search evidence: $\mathcal{X}_{I_{k}}$ Narrative assessment: 1. For each evidence cluster $\mathcal{C}_{k}^{\prime}$, the LLM compares the claim $T_{P}$ with the text component of $\mathcal{X}_{T_{k}}$ and $\mathcal{X}_{I_{k}}$. 2. The LLM conducts entity-level comparisons across five alignment dimensions: Location Named person Date Main topic Common objects 3. Based on alignment, with $\mathcal{X}_{T_{k}}$ and $\mathcal{X}_{I_{k}}$, each Narrative $\mathcal{C}_{k}^{\prime}$ is labeled as: Supports – entities align and no conflict Conflicts – misalignment in location, person, or date Irrelevant – lacks named entities or visually unrelated Reasoning with narratives: 4. LLM applies a decision strategy prioritizing reverse image search evidence: If any Narrative with valid $\mathcal{X}_{I_{k}}$ supports the claim → label as True If Narratives with valid $\mathcal{X}_{I_{k}}$ conflict or are irrelevant: Check Narratives with $\mathcal{X}_{T_{k}}$ If any Narratives with $\mathcal{X}_{T_{k}}$ supports → label as True If all Narratives with $\mathcal{X}_{T_{k}}$ conflict or are irrelevant → label as Misleading Explanation: 5. The LLM generates a textual justification grounded in the above alignment assessments.
Figure 3. Reasoning protocol used by the LLM (GPT4) to assess the veracity of a post $(T_{P},I_{P})$ using clustered direct ($\mathcal{X}_{T_{k}}$) and reverse image search ($\mathcal{X}_{I_{k}}$) evidence. Narrative assessment includes comparison and labeling, while reasoning prioritizes reverse image evidence to determine a final verdict.
4. Datasets
4.1. Current datasets
@@
Current misinformation datasets present several limitations that hinder the comprehensive analysis of real-world misinformation on social media. Many are synthetically generated(Liu et al., 2024b; Luo et al., 2021; Abdelnabi et al., 2022), which while useful for specific evaluations, often fail to capture the nuanced and evolving nature of online misinformation.
@@
In (Luo et al., 2021), some true samples used generic file images unrelated to the claim text, with their truth determined by the context in which news websites used these images. Evidence collected by (Abdelnabi et al., 2022) at the time aligned with this context, but many of these images have since been repurposed, causing misalignment in reverse image search evidence. Moreover, these datasets may not adequately reflect the linguistic complexities, contextual variations, and multimodal characteristics of real-world fake news.
@@
Thus it is important to focus on real-world fake news datasets as introduced in (Papadopoulos et al., 2024c; Tonglet et al., 2025). However, these datasets contain curated and sanitized evidence, which bypasses the significant challenge of retrieving relevant information from the vast and noisy web in practical scenarios. This is particularly problematic when dealing with divisive content, where different media outlets and individuals may present conflicting narratives and supporting evidence, a phenomenon often overlooked by traditional fact-checking models that tend to seek a singular truth. Consequently, there is a pressing need for datasets that better represent the complexities of real-world fake news, especially those surrounding topics where differing opinions and diverse sources contribute to a fragmented information landscape.
@@
MMbase(Liu et al., 2024b) - - - 0.76 - - -
AITR (Papadopoulos et al., 2024b) 0.52((0.55) 0.73(0.80) 0.52(0.49) 0.59 0.49 (0.52) 0.90 0.64
MUSE (Papadopoulos et al., 2024b) 0.53(0.57) 0.74( 0.80) 0.52(0.51) 0.58 0.49 0.89 0.63
RED-DOT(Papadopoulos et al., 2024a) - 0.76 - 0.60 0.46 0.88 0.61
COVE (Tonglet et al., 2025) - - - - 0.58 0.86 -
@@
Table 1. Accuracy comparison against state-of-the-art, across datasets. T/O : True versus OOC, T/F : True versus Fake, T/M: True versus Miscaptioned. For Verite T/F, we consider both OOC and Miscaptioned to be fake. We report the accuracy using the evidence that came with the datasets in Italics. For NewsCLIPpings we don’t collect evidence and use the evidence that came with the dataset. For DP, the numbers in parentheses indicate the accuracy only using the community notes as evidence.
5. Experiments and results
5.1. Comparison against state-of-the-art
@@
We focus our efforts on real-world fake news datasets where evidence retrieval and refinement can be more effectively applied to improve performance. Table 1 demonstrates that the binary Veracity prediction results of our zero-shot model CRAVE improve upon the state-of-the-art results on both real-world datasets Verite and 5PilsOOC, and the Synthetic dataset MMFakeBench for trained and zero-shot models alike. For NewsCLIPpings, due to the issues outlined in Sec. 4.1, we refrain from evidence collection or refinement, instead use the evidence provided (Abdelnabi et al., 2022). We apply an ablated version of CRAVE that uses only the Judgment step, achieving competitive results with the zero-shot model COVE, though not matching the performance of fully trained models.
5.2. Ablation studies
@@
Table 2 demonstrates the effectiveness of iterative refinement during text evidence $\mathcal{X}_{T}$ retrieval which leads to improved scores across datasets. Clustering of all evidence into distinct narratives leads to improved veracity prediction compared to taking the maximally similar evidence. Finally Agentic Retrieval of $\mathcal{X}_{T_{k}}$ based on $K$ individual narrative clusters leads to further improvements for all datasets except Verite, this could be attributed to the lack of novel evidence retrieval that is already not part of the evidence set.
@@
Table 2. Ablation Studies: Improvements due to text evidence refinement ( $\mathcal{X}_{T}$ Refine), Clustering before judgment(Cluster Judge) and Evidence refinement per cluster ( $\mathcal{X}_{T_{k}}$ Refine).
5.3. Ablation Studies on Clustering
For the veracity prediction task, we experiment with K-means, based on text embedding of the evidence, and prompt-based GPT 4 clustering on a subset (300 samples) of Verite. Row 1 in Table 3 denotes the initial veracity when taking the maximally similar evidence before clustering, and is thus the same for all methods. Row 2 denotes the performance when the judgment prompt uses the formed clusters using different clustering schemes.
We observe that K-means with $K=8$ provides the best performance, as it allows the LLM to more easily identify True Claims, which often form exact matches within a single cluster. However, this high cluster count complicates the refinement stage, as agents at $K=8$ perform individual searches for refinement. With $H=2$ searches, the total number of searches increases from 80 for $K=4$ to 160 for $K=8$, making the $K=8$ setup resource-intensive and unrealistic when applying to multiple datasets. As a result, we select $K=4$ for our design.
We also compare against a GPT-4 prompt-based clustering. While the initial clustering did not yield better judgments, GPT-4 often excluded irrelevant evidence from the clusters. As a result, in Row 3, we observe that using GPT-4 clusters for refining evidence led to the largest improvement in our ablated model, where we select the maximally similar evidence for veracity prediction.
Kmeans K=2 Kmeans K=4 Kmeans K=8 GPT4 K $\leq$ 4
Acc F1 Acc F1 Acc F1 Acc F1
CRAVE(max) before 0.74 0.73 0.74 0.73 0.74 0.73 0.74 0.73
CRAVE 0.79 0.78 0.80 0.79 0.83 0.82 0.79 0.78
CRAVE(max) after 0.73 0.72 0.75 0.74 0.75 0.73 0.77 0.76
Table 3. Effect of cluster Size K on veracity prediction: Increasing cluster size K improves veracity prediction but raises search costs; we fix K=4 for tractability.
Method $k$ 5PilsOOC Verite DP
Sil ↑ DB ↓ Sil ↑ DB ↓ Sil ↑ DB ↓
kmeans 2 0.29 1.31 0.25 1.46 0.29 1.28
4 0.32 1.19 0.28 1.26 0.30 1.18
8 0.35 0.78 0.31 0.81 0.33 0.78
spectral 2 0.25 1.55 0.21 1.70 0.24 1.54
4 0.19 1.47 0.18 1.50 0.19 1.46
8 0.13 1.18 0.11 1.20 0.11 1.20
Table 4. Cluster coherence metrics for different methods and values of $k$. For all datasets, k-means consistently outperforms spectral clustering across all values of $k$. Additionally, as $k$ increases, the Silhouette score generally improves, while the Davies-Bouldin index decreases, indicating better cluster coherence for both methods.

### v2

@@ III-C Reasoning: prompt-based comparison and explanation
Our Judgment Phase comprises comparison with the Claim leading to an interpretable explanation and coherent verdict. Our prompt, Figure 3, evaluates the veracity of a social media post $P=(T_{P},I_{P})$, consisting of a textual claim $T_{P}$ and an associated image $I_{P}$, by leveraging clustered evidence obtained from both direct and reverse image search. Each narrative cluster $\mathcal{C}_{k}^{\prime}$ contains evidence from text-based search ($\mathcal{X}_{T_{k}}$) and evidence from reverse search ($\mathcal{X}_{I_{k}}$).
III-C1 Narrative assessment
@@
The judgment process begins with narrative assessment, where the LLM compares the textual content of each evidence item in $\mathcal{C}_{k}^{\prime}$ against the claim text $T_{P}$. Unlike prior methods that treat context prediction as a standalone step [12], our method integrates contextual understanding directly into a unified prompting strategy. Using a 5W1H-inspired questioning framework—focusing on who, what, when, where, why, and how—the LLM evaluates alignment between the claim and evidence across five critical dimensions: named entities, dates, locations, key topics, and common objects. This entity-level comparison helps determine whether an evidence item supports the claim, conflicts with it, or lacks relevant content entirely. In real-world scenarios, each narrative cluster often includes a mixture of evidence with varying degrees of fine-grained support. Such variability introduces noise into the reasoning process, making it challenging to draw accurate conclusions if relying solely on the highest-ranked evidence within a narrative. To address this, we aggregate evidence holistically at the cluster level, treating each $\mathcal{C}_{k}^{\prime}$ as a coherent narrative. Rather than favoring individual pieces of evidence in isolation, the LLM considers the collective alignment pattern within each cluster. The output of this stage is a structured set of alignment labels for each narrative, guiding further reasoning.
III-C2 Reasoning with narratives
@@
Following narrative assessment, where each narrative cluster $\mathcal{C}_{k}^{\prime}$ is labeled based on entity-level alignment, the LLM proceeds to reason over these narratives to derive a final veracity decision. This reasoning process is grounded in the principle that narrative coherence, especially when supported by image provenance, is critical for verifying factual claims. Specifically, the LLM first identifies whether there exists any narrative cluster that both (i) contains valid reverse image search evidence ($\mathcal{X}_{I_{k}}$), and (ii) has been labeled as Supports during narrative assessment. If such a narrative is found, the post is immediately labeled as True. If no supporting narrative with valid reverse evidence is available, the model then checks for supporting narratives that contain only text-based evidence ($\mathcal{X}_{T_{k}}$). If at least one such narrative supports the claim, the label remains True. However, if all narratives—whether supported by image or text searches—are either labeled as Conflicts or Irrelevant, the post is ultimately labeled as Misleading. This structured approach ensures that multiple sources of evidence are considered in a tiered fashion, reducing the risk of false negatives or misclassifications due to noisy or incomplete top-ranked results. Finally, the LLM produces a natural language explanation grounded in the alignment rationale, clearly articulating the justification behind the final verdict and enhancing the transparency and interpretability of the model’s reasoning. To ensure transparency and facilitate the reproduction of our results, the code and data supporting this study will be made publicly available upon acceptance.
IV Datasets
IV-A Current datasets
@@
Current misinformation datasets present several limitations [1] that hinder the comprehensive analysis of real-world misinformation on social media. Many are synthetically generated [18, 10, 11], which, while useful for specific evaluations, often fail to capture the nuanced and evolving nature of online misinformation.
@@
In [10], some true samples used generic file images unrelated to the claim text, with their truth determined by the context in which news websites used these images. Evidence collected by [11] at the time aligned with this context, but many of these images have since been repurposed, causing misalignment in reverse image search evidence. Moreover, these datasets may not adequately reflect the linguistic complexities, contextual variations, and multimodal characteristics of real-world fake news.
@@
Thus it is important to focus on real-world fake news datasets as introduced in [9, 12]. However, these datasets contain curated and sanitized evidence, which bypasses the significant challenge of retrieving relevant information from the vast and noisy web in practical scenarios. This is particularly problematic when dealing with divisive content, where different media outlets and individuals may present conflicting narratives and supporting evidence, a phenomenon often overlooked by traditional fact-checking models that tend to seek a singular truth. Consequently, there is a pressing need for datasets that better represent the complexities of real-world fake news, especially those surrounding topics where differing opinions and diverse sources contribute to a fragmented information landscape. To address these gaps, we construct a dataset that captures the complexity of online discourse—particularly the challenge of verifying claims amid multiple, and sometimes conflicting, perspectives and evidence.
Characteristic Count
Total tweets 608
Labels: True 359
Labels: Misleading 225
Labels: Conflicted 8
Average notes per tweet 2.54
Average ratings per note 138
Notes labeled as ”helpful” 0.59
Notes labeled as ”unhelpful” 0.37
TABLE I: Statistical overview of the DP dataset.
IV-B Divisive Posts (DP) evaluation benchmark
We constructed our multimodal dataset from X (formerly Twitter) Community Notes, a collaborative fact-checking system where contributors identify potentially misleading information. Our collection process followed a systematic approach to ensure quality and relevance. We systematically curated our dataset by first collecting daily Community Notes and their corresponding ratings from X’s transparency initiative. We then performed frequency analysis on note records, prioritizing tweets with the highest number of notes to target content with high engagement and potential controversy. For each selected tweet, we integrated ratings by mapping and filtering to the top 5 rated notes, ensuring quality and diverse perspectives. Using the X API, we retrieved the full tweet content, including text and media, to support multimodal analysis. Only tweets containing both text and at least one image were retained. Finally, human annotators manually labeled each tweet, referencing external sources to establish ground truth veracity. The resulting dataset, details in Table I, provides a benchmark for evaluating multimodal misinformation detection models, as it captures real-world instances of potentially misleading content along with community-generated assessments. While Community Notes data is publicly available, we ensured ethical usage by removing personal identifiers, focusing only on public account content, and fully complying with X’s developer terms of service.
V Experiments and results
@@
MMbase [18] - - - 0.76 - - -
AITR [8] 0.52((0.55) 0.73(0.80) 0.52(0.49) 0.59 0.49 (0.52) 0.90 0.64
MUSE [8] 0.53(0.57) 0.74( 0.80) 0.52(0.51) 0.58 0.49 0.89 0.63
RED-DOT [7] - 0.76 - 0.60 0.46 0.88 0.61
COVE [12] - - - - 0.58 0.86 -
@@
TABLE II: Accuracy comparison against state-of-the-art methods, across datasets. T/O : True versus Out-Of-Context(OOC), T/F : True versus Fake, T/M: True versus Miscaptioned. For Verite T/F, we consider both OOC and Miscaptioned to be fake. For Verite and 5PilsOOC we report the accuracy using the evidence that came with the datasets in italics. For NewsCLIPpings we don’t collect evidence and use the evidence that came with the dataset. For DP, the numbers in parentheses indicate the accuracy only using the community notes as evidence.
V-A Comparison against state-of-the-art methods
@@
We focus our efforts on real-world fake news datasets where evidence retrieval and refinement can be more effectively applied to improve performance. Table II demonstrates that the binary veracity prediction results of our zero-shot model CRAVE improve upon the state-of-the-art results on both real-world datasets Verite and 5PilsOOC, as well as the Synthetic dataset MMFakeBench. For NewsCLIPpings, due to the issues outlined in subsection IV-A, we refrain from evidence collection or refinement, instead use the evidence provided [11]. We apply an ablated version of CRAVE that uses only the LLM reasoning step, achieving competitive results with the zero-shot model COVE, though not matching the performance of fully trained models. The table also shows that the performance of these trained models degrade when dealing with our noisy web-scraped evidence data versus the curated and cleaned evidence data that came with the datasets(denoted in parentheses). While these methods use different pipelines for evidence collection, we use a single evidence collection pipeline, refined through clustering and filtering. For DP, we also test against the Community notes data to validate the superior quality of our collected evidence.
V-B Baseline performance using MLLMs
Table III compares Multimodal Large Language Models (MLLMs) with our CRAVE framework on the binary veracity prediction task. For both InstructBLIP [27] and LLAVA [44], we use Vicuna-13B [28] as the underlying language model. All models are evaluated on the same input: the claim and the retrieved evidence. However, performance varies significantly due to how they process and reason over noisy, multimodal contexts.
Method Verite MMFakebench 5PilsOOC NewsCLIPpings DP
Acc Acc Acc Acc Acc
InstructBLIP (C) 0.50 0.50 0.55 0.55 0.54
InstructBLIP (V) 0.63 0.62 0.50 0.50 0.40
InstructBLIP (T) 0.56 0.50 0.50 0.47 0.43
LLAVA (C) 0.66 0.68 0.53 0.54 0.40
LLAVA (T) 0.68 0.71 0.60 0.69 0.49
GPT 4 (V) 0.78 [22] 0.75 [18] - - -
CRAVE 0.82 0.78 0.81 0.85 0.73
TABLE III: Comparison against MLLMs across datasets. C denotes models using only the Claim. V denotes the version using both images and text as evidence, and T denotes versions using only the text component of the evidence.
Additionally, unlike MLLMs—whose performance may be skewed by biases from pretraining on large-scale web data—CRAVE is less influenced by domain priors or cultural biases. This is particularly advantageous in datasets like 5PilsOOC and DP, where MLLMs struggle with claims from underrepresented regions or topics. Finally, standard prompting – asking a powerful model to fact-check an image/text pair – often fails to deal with conflicting noisy evidence. CRAVE’s zero-shot approach and reliance on structured narratives lead to more robust decision-making, regardless of cultural or regional context.
V-C Ablation study
To validate our design choices and quantify the contribution of each module, we conduct a comprehensive ablation study. We first dissect the impact of our framework’s core architectural components, such as iterative refinement and narrative clustering. We assess the system’s sensitivity to external factors, evaluating various backbone LLMs and evidence modalities to understand their influence on overall performance. Finally, we then delve into the clustering process itself, analyzing the optimal cluster count and the choice of method.
V-C1 Study 1 - Role of components
@@
Table IV demonstrates the effectiveness of iterative refinement during text evidence $\mathcal{X}_{T}$ retrieval which leads to improved scores across datasets. Clustering of all evidence into distinct narratives leads to improved veracity prediction compared to taking the maximally similar evidence. Finally Agentic Retrieval of $\mathcal{X}_{T_{k}}$ based on $K$ individual narrative clusters leads to further improvements for all datasets except Verite; this could be attributed to the lack of novel evidence retrieval that is already not part of the evidence set.
@@
TABLE IV: Role of components: Improvements due to text evidence refinement ( $\mathcal{X}_{T}$ Refine), Clustering before judgment(Cluster Judge) and Evidence refinement per cluster ( $\mathcal{X}_{T_{k}}$ Refine).
Figure 4: Dynamic Clustering performance varies with dataset complexity, as higher similarity thresholds fragment evidence into more clusters. At peak entropy—where clustering uncertainty is highest—samples span 4 clusters on average. Since the optimal threshold is claim-dependent and cannot be determined a priori, we adopt a static clustering with K=4.
V-C2 Study 2 - Backbone LLM and evidence stream
Backbone LLM for Judgment Verite 5PilsOOC DP
GPT-4o 0.82 0.81 0.73
Gemini-2.0-Flash 0.77 0.83 0.74
Llama-3.1-8B-Instruct 0.77 0.77 0.67
Llama-3.1-8B-Instruct (I) 0.75 0.75 0.69
Llama-3.1-8B-Instruct (T) 0.72 0.73 0.66
TABLE V: Backbone LLM and evidence Stream. Commercial LLMs excel due to superior reasoning; reverse image search(I) outperforms text-only(T), with T+I combination being optimal
In Table V we study various backbone LLMs for veracity prediction accuracy. Commercial models perform on par, with a notable boost on the complex DP dataset, likely due to superior reasoning. However, even with open models like Llama we improve upon the current state of the art. Unsurprisingly image provenance through Reverse image search proves to be a stronger signal that text search (T) alone. While combining them (T+I) usually leads to better performance, on the divisive DP dataset it often gets overwhelmed by conflicting narratives from the text evidence.
V-C3 Study 3 - Cluster size
Our analysis of Dynamic(Agglomerative) Clustering performance, Figure 4, reveals that the optimal configuration is intrinsically tied to the complexity of the underlying dataset. We identify the optimal similarity threshold, by locating the peak of the clustering entropy—a point representing maximum algorithmic uncertainty and a critical transitional state. For datasets like 5PILSOOC and Verite, curated from fact-checking corpora, this heuristic aligns well with peak classification accuracy. 5PILSOOC, with its verbatim claims, achieves effective classification at a low threshold (0.70) requiring an average of only 2.7 clusters. Verite, featuring paraphrased claims, introduces greater semantic ambiguity, thus demanding a higher threshold (0.80) and a larger average cluster count (K=4.5) to resolve these nuances.
In stark contrast, our DP dataset, derived from informal Community Notes, presents a more complex challenge. Here, peak accuracy ($73.2\%$) is achieved only at a stringent threshold of 0.90, well after the point of peak entropy (which occurs at 0.85). This divergence demonstrates that for nuanced real-world fake claims, the system must be pushed beyond its state of maximum uncertainty, requiring an average of 6.4 clusters to isolate the correct classifying signal from a large set of distractors.
This variance underscores a critical insight: an optimal dynamic clustering threshold is claim-dependent and cannot be determined a priori for a general-purpose system. Lacking a reliable method to predict claim complexity for an adaptive threshold, we adopt a static clustering with K=4, a value empirically grounded as the average number of clusters per sample (4.0) observed at peak entropy across our datasets.
Method 5PilsOOC Verite DP
Sil ↑ DB ↓ Acc ↑ Sil ↑ DB ↓ Acc ↑ Sil ↑ DB ↓ Acc ↑
Kmeans 0.32 1.19 0.81 0.28 1.26 0.80 0.30 1.18 0.70
GPT 4o-mini 0.35 1.08 0.78 0.33 1.12 0.79 0.31 1.21 0.67
Spectral 0.19 1.47 0.74 0.18 1.50 0.72 0.19 1.46 0.66
TABLE VI: Cluster coherence metrics for different method. For all datasets, k-means consistently outperforms spectral clustering.
V-C4 Study 4 - Cluster method
The clustering performance metrics, in Table VI, reveal key insights into the effectiveness of different methods across random subsets (300 samples each) of the real-world datasets 5PilsOOC, Verite, and DP. The evaluation metrics include the average silhouette score (Sil) [45] (indicating cluster cohesion and separation, with higher values being better) and the Davies-Bouldin(DB) [46] index (where lower values indicate better-defined clusters) along with the veracity accuracy. We also compare against a GPT 4o-mini prompt-based clustering. While the cluster coherence was competitive, if not better at times, GPT 4o-mini often excluded outlier evidence from the clusters.